# RACNN: Residual Attention Convolutional Neural Network for Near-Field Channel Estimation in 6G Wireless Communications

### v1

@@ Abstract
Near-field channel estimation is a fundamental challenge in sixth-generation (6G) wireless communication, where extremely large antenna arrays (ELAA) enable near-field communication (NFC) but introduce significant signal processing complexity. Traditional model-based methods suffer from high computational costs and limited scalability in large-scale ELAA systems, while existing learning-based approaches often lack robustness across diverse channel conditions. To overcome these limitations, we propose the Residual Attention Convolutional Neural Network (RACNN), which integrates convolutional layers with self-attention mechanisms to enhance feature extraction by focusing on key regions within the CNN feature maps. Experimental results show that RACNN outperforms both traditional and learning-based methods, including XLCNet, across various scenarios, particularly in mixed far-field and near-field conditions. Notably, in these challenging settings, RACNN achieves a normalized mean square error (NMSE) of $4.8*10^{-3}$ at an SNR of 20dB, making it a promising solution for near-field channel estimation in 6G.
@@ 1 Introduction
Channel estimation is an essential problem in wireless communication systems and many studies have been proposed to address this problem in far-field scenarios [3]. These studies can be divided into two groups: (i) traditional model-based methods, such as least squares (LS) and minimum mean square error (MMSE) [8], and (ii) learning-based methods, which have recently gained attention for leveraging machine learning techniques in channel estimation. Methods in the former group are simple to implement in practice. However, their estimation accuracy is affected by noise and interference. Moreover, when the number of ELAA elements grows larger, the matrix inversion required in these methods is computationally challenging. The latter group is further divided into two approaches, i.e., model-driven methods and data-driven methods. The model-driven approach focuses on optimizing the model based on mathematical solutions. For example, in [14], the authors presented LISTA and SDL-LISTA networks, which are neural networks based on the iterative shrinkage and thresholding algorithm (ISTA). In [9], the authors proposed AMP-NET based on the principle of approximate message passing (AMP) for joint activity detection and channel estimation. These methods achieve higher estimation accuracy compared to the original methods with a small number of learning parameters. However, a neural network with a small number of learning parameters also has its own drawbacks. They perform well only under specific channel conditions that match their training data, such as the number of far-field paths, near-field paths, and noise distribution. The data-driven approach can mitigate this limitation by training neural networks with a large number of learning parameters. This enables the networks to learn from a wide range of channel conditions. To illustrate, in [2], the authors proposed XLCNet for channel estimation in NFC-6G systems. Specifically, XLCNet transforms noisy channel vectors into 2-channel images and then uses a convolutional neural network (CNN) to denoise them. Notably, this network supports not only far-field and near-field scenarios but also mixed far-field/near-field scenarios. Moreover, CNN is a widely studied architecture that has been utilized for many years. With recent advancements in learning-based techniques for image processing and natural language processing, the underlying concept of XLCNet has the potential for further enhancement.
@@
Inspired by [2] and a recent advancement in deep learning (DL) known as the attention mechanism [11], we propose a novel network named the residual attention convolutional neural network (RACNN) for NFC-6G wireless communication. [^1] Simulation codes: https://github.com/DoHaiSon/RACNN. RACNN inherits the idea of estimating noise through a neural network and subtracting the estimated noise from the input at the final stage. RACNN follows the convolutional block structure of XLCNet and incorporates self-attention blocks to enhance the overall performance of noise estimation. Our numerical results show that RACNN outperforms traditional model-based methods (LS and MMSE) and XLCNet in various scenarios, e.g., (i) the testing dataset matches the number of far-field and near-field paths in the training dataset; (ii) only far-/near-field paths exist; and (iii) mixed far-field/near-field paths. The rest of this paper is organized as follows. Section 2 introduces the Rayleigh distance and the channel model in NFC-6G. Section 3 presents our proposed RACNN network for channel estimation. Experiment setups, evaluation methodology, results, and performance analysis are discussed in details in Section 4. Finally, we conclude the paper in Section 5.
@@ 2.2 Channel Models
We consider an NFC-6G wireless channel with an ELAA consisting of $M$ elements, serving $K$ user equipments. Following the article in [13], the downlink far-field channel is modeled as
$\mathbf{Y}=\mathbf{H}\mathbf{P}^{\frac{1}{2}}\mathbf{X}+\mathbf{Z},$ (2)
where $\mathbf{H}\in\mathbb{C}^{M\times K}$ represents the channel matrix, $\mathbf{P}=\operatorname{diag}\left(\mathit{P}_{1},\mathit{P}_{2},\dots,\mathit{P}_{K}\right)$ is the signal amplitude, $\mathbf{X}\in\mathbb{C}^{K\times K}$ is the matrix of data symbols, and $\mathbf{Z}\in\mathbb{C}^{M\times K}$ is an additive white Gaussian noise matrix. We also assume that signal transmission is lossless and signal amplitudes across all antennas are identical, i.e., $P_{k}=P~{}\forall k$. In NFC-6G systems, the channel matrix $\mathbf{H}$ can be expressed in a unified form for both far-field and near-field cases as follows:
@@
Here, each channel vector $\mathbf{a}(\phi_{l})$ is expressed as
@@
where $d=\frac{\lambda}{2}$ denotes the antenna spacing within the ELAA and $\phi_{l}=\left[-\frac{\pi}{2};\frac{\pi}{2}\right]$ be the steering angle of the $l^{\text{th}}$ path. On the other hand, the steering matrix $\mathbf{A}_{\text{near}}$ in near-field cases is parameterized by $\phi_{l}$ and $r_{l}$ as follows
@@
where $r_{l}$ represents the distance from the $l^{\text{th}}$-antenna element to the center of the antenna array, while $r_{l,m}=\sqrt{r^{2}_{l}+\delta^{2}_{m}d^{2}-2r_{l}\delta_{m}d\sin(\phi_{l})}$ is the distance between the $l^{\text{th}}$-scatterer and the $n^{\text{th}}$-antenna element at the base station with $\delta_{m}=\frac{2n-N-1}{2}$ and $m=1,2,\cdots,M$.
@@
We refer the readers to [13] for further details on these channel models.
@@ 3 Proposed Residual Attention CNN Network
Here, ``"``"`$\star$"` ` ⋆ " denotes convolution operator, the weight $\mathbf{W}$ is a kernel matrix, $\mathbf{W}\in\mathbb{R}^{r\times r}$ (e.g., $r=3$), and $\texttt{bias}(C_{y})$ represents the bias to alter the activation function. $C_{x}$ and $C_{y}$ are the channel numbers of input $\mathbf{X}$ and output $\mathbf{Y}$, respectively. In our setting, we use $64$ filters in each conv2d block and set the value of $C_{x}$ to $2$.
To reduce overfitting during the training process, the output from certain conv2D blocks passes through a rectified linear unit (ReLU) activation function [5] to introduce nonlinearity. Batch normalization (BN) [6] is also employed to adjust and scale activations, helping reduce the covariate shifting. Specifically, the BN in our network is defined by
${\mathbf{Y}_{\text{BN}}}=\gamma{\mathbf{Y}_{\text{conv2d}}}+\beta,$ (11)
where $\gamma$ and $\beta$ are the scale and shift parameters, ${\hat{\mathbf{Y}}_{\text{conv2d}}}$ is normalized from the input ${\mathbf{Y}_{\text{conv2d}}}$ ( output of the conv2d layer) as ${\hat{\mathbf{Y}}_{\text{conv2d}}}=\frac{{\mathbf{Y}_{\text{conv2d}}}-\mu_{B}}{\sqrt{\sigma^{2}_{B}+\epsilon}}$, with $\mu_{B}$ and $\sigma^{2}_{B}$ is the mean and variance of ${\mathbf{Y}_{\text{conv2d}}}$.
To support the extraction stage in each conv2D, a Residual Attention (RA)block is paired with the network to introduce the attention mechanism. An RA block has two mechanisms: (i) the data attention function refines the features learned by the network, and (ii) the residual connection function helps maintain effective network learning as data progresses deeper [12]. Particularly in our RACNN network, we use the self-attention mechanism based on [11]. This layer generates an attention map that highlights the important regions or features of the input using its own weights. Accordingly, RA block helps each output from conv2D blocks retain the original information while incorporating the refined features from the attention mechanism. Fig. 1(c) illustrates the architecture of a Residual Attention block, which follows the following sequence:
@@ 4.1 Experiment Setup
Wavelength $\lambda=0.01$ No. interations 200
@@
$\text{NMSE}=\frac{\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},\qquad\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2},$ (16)
where $y_{i}$ is the observed value, $\hat{y_{i}}$ is the estimated value, $\bar{y}$ is the mean of the observed values, and $n$ is the total number of observations.
@@ 4.2.2 Far-Field and Near-Field Channel Estimation:
In this experiment, we validate the performance of RACNN on near-field channel and far-field channel independently. This is done by modifying the BS transmit path $L_{f}$ and $L_{n}$ of each electromagnetic field in a range of constant (3, 5, and 7). The experimental results are shown in Fig. 3 and Fig. 4. They indicate that RACNN achieves the best estimation accuracy performance in most testing cases. At low SNRs ($\leq$4dB), MMSE performs better, particularly with lower multipath values $L_{f}=3$ and $L_{n}=3$. However, when the multipath increases, the performance of RACNN begins to improve significantly, as compared to others. Fig. 5 shows the detailed results, investigating channel configurations. We can observe that all methods perform poorly at lower SNR levels. As the SNR increases, a noticeable difference in performance emerges. Looking at the far-field channel, changing in channel configuration have a small impact on impact on XLCNet’s results. In contrast, with RACNN, each line becomes progressively more differentiated as the SNR increases. This behavior differs when the same test is conducted for the near-field channel, where both RACNN and XLCNet exhibit stable performance across all scenarios despite the increasing SNR.
@@ 4.2.3 Hybrid-Field Channel Estimation:
Similar to the experiments above, we also apply the same validation approach for the hybrid-field channel by mixing different $L_{f}$ and $L_{n}$ settings. The test begins with the lowest multipath configuration at $L_{f}=3$ and $L_{n}=3$, then gradually increases far-field and near-field paths to different values (e.g., 6 or 9) to investigate the impact of system configuration changes on RACNN and its competitors. The experimental results, shown in Fig. 8, indicate that RACNN maintains its position of the best performance on hybrid-field channel except at SNR of 0dB. Consistent with the results of previous experiments, RACNN’s performance improves as SNR increases, outperforming XLCNet at high SNRs. From Fig. 8, we also observe that RACNN is a good option for hybrid-field channel estimation, achieving good performance, especially at high SNRs with fewer multipaths. The difference in RACNN performance is expressed prominently at the high SNR range, especially in the range of 16 to 20dB, where the error between the best and worst result is around $2*10^{-3}$. From here, we expect a potential where RACNN capability could be improved from the current result with a few tweaks in model architecture. Despite the performance differences, an increase in channel paths does not directly lead to noise reduction, as performance ranking remains inconsistent across SNR levels. This is evident in Fig. 8, where we observed the advantage of increased transmitting multipath at lower SNR. However, as SNR rises, channel configurations with higher multipaths begin to pose a drawback. This property applies to both far-field and near-field channels.
@@ 5 Conclusion
In this paper, we proposed the Residual Attention Convolutional Neural Network for near-field channel estimation in 6G wireless communication systems. RACNN integrates convolutional layers with self-attention mechanisms to enhance feature extraction during donoising process. Through various experiments, RACNN consistently outperformed other estimators, i.e., LS, MMSE, and XLCNet. More details, in scenarios with different far-field and near-field path settings, RACNN maintained the best performance across all cases, except at 0dB SNR where MMSE had a slight advantage. The experimental data also indicated that RACNN’s performance improved significantly as the number of near-field and far-field paths reduced. Overall, the findings suggest that RACNN is a promising solution for the complex signal processing demands of ELAA in 6G systems, offering robustness and improved accuracy in near-field channel estimation.

### v3

@@ Abstract
Near-field channel estimation is a fundamental challenge in the sixth-generation (6G) wireless communication, where extremely large antenna arrays (ELAA) enable near-field communication (NFC) but introduce significant signal processing complexity. Traditional model-based methods suffer from high computational costs and limited scalability in large-scale ELAA systems, while existing learning-based approaches often lack robustness across diverse channel conditions. To overcome these limitations, we propose the Residual Attention Convolutional Neural Network (RACNN), which integrates convolutional layers with self-attention mechanisms to enhance feature extraction by focusing on key regions within the CNN feature maps. Experimental results show that RACNN outperforms both traditional and learning-based methods, including XLCNet, across various scenarios, particularly in mixed far-field and near-field conditions. Notably, in these challenging settings, RACNN achieves a normalized mean square error (NMSE) of $4.8*10^{-3}$ at an SNR of 20dB, making it a promising solution for near-field channel estimation in 6G.
@@ 1 Introduction
Channel estimation is an essential problem in wireless communication systems and many studies have been proposed to address this problem in far-field scenarios [3]. These studies can be divided into two groups: (i) traditional model-based methods, such as least squares (LS) and minimum mean square error (MMSE) [8], and (ii) learning-based methods, which have recently gained attention for leveraging machine learning techniques in channel estimation. Methods in the former group are simple to implement in practice. However, their estimation accuracy is affected by noise and interference. Moreover, when the number of ELAA elements grows larger, the matrix inversion required in these methods is computationally challenging. The latter group is further divided into two approaches, i.e., model-driven methods and data-driven methods. The model-driven approach focuses on optimizing the model based on mathematical solutions. For example, in [14], the authors presented learning ISTA (LISTA) and sparsifying dictionary learning-LISTA (SDL-LISTA) networks, which are neural networks based on the iterative shrinkage and thresholding algorithm (ISTA). In [9], the authors proposed AMP-NET based on the principle of approximate message passing (AMP) for joint activity detection and channel estimation. These methods achieve higher estimation accuracy compared to the original methods with a small number of learning parameters. However, a neural network with a small number of learning parameters also has its own drawbacks. They perform well only under specific channel conditions that match their training data, such as the number of far-field paths, near-field paths, and noise distribution. The data-driven approach can mitigate this limitation by training neural networks with a large number of learning parameters. This enables the networks to learn from a wide range of channel conditions. To illustrate, in [2], the authors proposed an extremely large-scale MIMO channel network (XLCNet) for channel estimation in NFC-6G systems. Specifically, XLCNet transforms noisy channel vectors into 2-channel images and then uses a convolutional neural network (CNN) to denoise them. Notably, this network supports not only far-field and near-field scenarios but also mixed far-field/near-field scenarios. Moreover, CNN is a widely studied architecture that has been utilized for many years. With recent advancements in learning-based techniques for image processing and natural language processing, the underlying concept of XLCNet has the potential for further enhancement.
@@
Inspired by [2] and a recent advancement in deep learning (DL) known as the attention mechanism [11], we propose a novel network named the RACNN for NFC-6G wireless communication. [^1] Simulation codes: https://github.com/DoHaiSon/RACNN. RACNN inherits the idea of estimating noise through a neural network and subtracting the estimated noise from the input at the final stage. RACNN follows the convolutional block structure of XLCNet and incorporates self-attention blocks to enhance the overall performance of noise estimation. Our numerical results show that RACNN outperforms traditional model-based methods (LS and MMSE) and XLCNet in various scenarios, e.g., (i) matched training and testing configurations in terms of far-/near-field paths; (ii) testing with only far-field or only near-field paths; and (iii) mismatched mixtures of far-field and near-field components.
@@ 2.2 Channel Models
We consider an NFC-6G wireless channel with an ELAA consisting of $M$ elements, serving $K$ UEs. Following the article in [2], the downlink far-field channel is modeled as
$\mathbf{Y}=\mathbf{H}\mathbf{P}^{\frac{1}{2}}\mathbf{S}+\mathbf{Z},$ (2)
where $\mathbf{H}\in\mathbb{C}^{M\times K}$ represents the channel matrix, $\mathbf{P}=\operatorname{diag}\left(\mathit{P}_{1},\mathit{P}_{2},\dots,\mathit{P}_{K}\right)$ is the transmit power matrix, $\mathbf{S}\in\mathbb{C}^{K\times K}$ is the transmit pilot matrix, and $\mathbf{Z}~{}\in~{}\mathbb{C}^{M\times K}$ is an additive white Gaussian noise matrix. We also assume that signal transmission is lossless and transmit powers across all antennas are identical, i.e., $P_{k}=P~{}\forall k$. In NFC-6G systems, the channel matrix $\mathbf{H}$ can be expressed in a unified form for both far-field and near-field cases [13] as follows:
@@
Here, consider a uniform linear array (ULA) at the BS, where each channel vector $\mathbf{a}(\phi_{l})$ is expressed as
@@
where $d=\frac{\lambda}{2}$ denotes the antenna spacing and $\phi_{l}=\left[-\frac{\pi}{2};\frac{\pi}{2}\right]$ be the steering angle of the $l$-th path. On the other hand, the steering matrix $\mathbf{A}_{\text{near}}$ in near-field case is parameterized by $\phi_{l}$ and $r_{l}$ as follows
@@
where $r_{l}$ represents the distance from the $l$-th scatter to the center of the antenna array, while $r_{l,m}=\sqrt{r^{2}_{l}+\delta^{2}_{m}d^{2}-2r_{l}\delta_{m}d\sin(\phi_{l})}$ is the distance between the $l$-th scatter and the $m$-th antenna element at the base station (BS) with $\delta_{m}=\frac{2m-M-1}{2}$ and $m=1,2,\cdots,M$.
@@
where $L_{o}$ and $L-L_{o}$ represent the number of far-field paths and near-field paths, respectively. For clarity, we introduce the notation $L_{f}=L_{o}$ and $L_{n}=L-L_{o}$ in the following sections of the paper. We refer the readers to [13] for further details on these channel models.
@@ 3 Proposed Residual Attention CNN Network
Here, ``"``"`$\star$"` ` ⋆ " denotes the 2D convolution operator, $\mathbf{W}(C_{y},k)\in\mathbb{R}^{r\times r}$ (e.g., $r=3$) is the convolution kernel connecting $k$-th input channel to $C_{y}$-th output channel, and $\texttt{bias}(C_{y})\in\mathbb{R}$ is the associated bias term. In our implementation, the first convolutional layer takes $C_{x}=2$ as input, and the number of output channels is set to $C_{y}=64$. In subsequent layers, the number of input and output channels remains fixed at $C_{x}=C_{y}=64$. At the final layer, the output channels are reduced to $C_{y}=2$, yielding an estimated noise that shares the same shape as the input $\mathbf{X}\in\mathbb{R}^{2\times N_{x}\times N_{y}}$.
To reduce overfitting during the training process, the output from certain conv2D blocks passes through a ReLU activation function [5] to introduce nonlinearity. BN [6] is also employed to adjust and scale activations, helping reduce the covariate shift. Specifically, the BN in our network is defined by
${\mathbf{Y}_{\text{BN}}}=\gamma\hat{\mathbf{Y}}_{\text{conv2d}}+\beta,$ (11)
where $\gamma$ and $\beta$ are the scale and shift parameters, ${\hat{\mathbf{Y}}_{\text{conv2d}}}$ is normalized from the input ${\mathbf{Y}_{\text{conv2d}}}$ (i.e., output of the conv2d layer) as ${\hat{\mathbf{Y}}_{\text{conv2d}}}=\frac{{\mathbf{Y}_{\text{conv2d}}}-\mu_{B}}{\sqrt{\sigma^{2}_{B}+\epsilon}}$, with $\mu_{B}$ and $\sigma^{2}_{B}$ is the mean and variance of ${\mathbf{Y}_{\text{conv2d}}}$.
To support the extraction stage in each conv2D, an RA block is paired with the network to introduce the attention mechanism. The RA block has two mechanisms: (i) the data attention function refines the features learned by the network, and (ii) the residual connection function helps maintain effective network learning as data progresses deeper [12]. Particularly in our RACNN network, we use the self-attention mechanism based on [11]. This layer generates an attention map that highlights the important regions or features of the input using its own weights. Accordingly, RA block helps each output from conv2D blocks retain the original information while incorporating the refined features from the attention mechanism. Fig. 1(c) illustrates the architecture of an RA block, which follows the following sequence:
@@ 4.1 Experiment Setup
Wavelength $\lambda=0.01$ No. iterations 200
@@
Near-field distance $r_{l}=\mathcal{U}(10;80)$ Training SNR levels $10$ and $15$dB
@@
$\text{NMSE}=\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}|y_{i}|^{2}},\qquad\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2},$ (16)
where $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and estimated values, respectively, and $n$ is the total number of observations.
@@ 4.2.2 Far-Field and Near-Field Channel Estimation:
In this experiment, we validate the performance of RACNN on near-field channel and far-field channel independently. This is done by modifying the BS transmit path $L_{f}$ and $L_{n}$ of each electromagnetic field in a range of constant (3, 5, and 7). The experimental results are shown in Fig. 3 and Fig. 4. They indicate that RACNN achieves the best estimation accuracy performance in most testing cases. At low SNRs ($\leq$4dB), MMSE performs better, particularly with lower multipath values $L_{f}=3$ and $L_{n}=3$. However, when the number of multipaths increases, the performance of RACNN begins to improve significantly, as compared to others. Fig. 5 shows the detailed results, investigating channel configurations. We can observe that all methods perform poorly at lower SNR levels. As the SNR increases, a noticeable difference in performance emerges. Looking at the far-field channel, changes in channel configuration have a small impact on XLCNet’s results. In contrast, with RACNN, each line becomes progressively more differentiated as the SNR increases. This behavior differs when the same test is conducted for the near-field channel, where both RACNN and XLCNet exhibit stable performance across all scenarios despite the increasing SNR.
@@ 4.2.3 Hybrid-Field Channel Estimation:
Similar to the experiments above, we also apply the same validation approach for the hybrid-field channel by mixing different $L_{f}$ and $L_{n}$ settings. The test begins with the lowest multipath configurations at $L_{f}=3$ and $L_{n}=3$, then gradually increases far-field and near-field paths to different values (e.g., 6 or 9) to investigate the impact of system configuration changes on RACNN and its competitors. The experimental results, shown in Fig. 8, indicate that RACNN maintains its position of the best performance on hybrid-field channel except at SNR of 0dB. Consistent with the results of previous experiments, RACNN’s performance improves as SNR increases, outperforming XLCNet at high SNRs. From Fig. 8, we also observe that RACNN is a good option for hybrid-field channel estimation, achieving good performance, especially at high SNRs with fewer multipaths. The difference in RACNN performance is expressed prominently at the high SNR range, especially in the range of 16 to 20dB, where the gap between the best and worst results reaches approximately $2*10^{-3}$. From here, we expect a potential where RACNN capability could be improved from the current result with a few tweaks in model architecture. Despite the performance differences, an increase in channel paths does not directly lead to noise reduction, as performance ranking remains inconsistent across SNR levels. This is evident in Fig. 8, where we observed the advantage of increased transmitting multipaths at lower SNR. However, as SNR rises, channel configurations with higher multipaths begin to pose a drawback. This property applies to both far-field and near-field channels.
@@ 5 Conclusion
In this paper, we proposed the Residual Attention Convolutional Neural Network for near-field channel estimation in 6G wireless communication systems. RACNN integrates convolutional layers with self-attention mechanisms to enhance feature extraction during denoising process. Through various experiments, RACNN consistently outperformed other estimators, i.e., LS, MMSE, and XLCNet. More details, in scenarios with different far-field and near-field path settings, RACNN maintained the best performance across all cases, except at 0dB SNR where MMSE had a slight advantage. The experimental data also indicated that RACNN’s performance improved significantly as the number of near-field and far-field paths reduced. Overall, the findings suggest that RACNN is a promising solution for the complex signal processing demands of ELAA in 6G systems, offering robustness and improved accuracy in near-field channel estimation.
5.0.1 Acknowledgments
This research was funded by the research project QG (QG.25.08) of Vietnam National University Hanoi. Correspondence: Tran Thi Thuy Quynh (quynhttt@vnu.edu.vn)
5.0.2 Disclosure of Interests
The authors have no competing interests.